### v1

# Growing a Twig to Accelerate Large Vision-Language Models

Abstract
Large vision-language models (VLMs) have demonstrated remarkable capabilities in open-world multimodal understanding, yet their high computational overheads pose great challenges for practical deployment. Some recent works have proposed methods to accelerate VLMs by pruning redundant visual tokens guided by the attention maps of VLM’s early layers. Despite the success of these token pruning methods, they still suffer from two major shortcomings: (i) considerable accuracy drop due to insensitive attention signals in early layers, and (ii) limited speedup when generating long responses (e.g., 30 tokens). To address the limitations above, we present TwigVLM—a simple and general architecture by “growing” a lightweight twig upon an early layer of the base VLM. Compared with most existing VLM acceleration methods purely based on visual token pruning, our TwigVLM not only achieves better accuracy retention by employing a twig-guided token pruning (TTP) strategy, but also yields higher generation speed by utilizing a self-speculative decoding (SSD) strategy. Taking LLaVA-1.5-7B as the base VLM, experimental results show that TwigVLM preserves 96% of the original performance after pruning 88.9% of visual tokens and achieves 154% speedup in generating long responses, delivering significantly better performance in terms of both accuracy and speed over the state-of-the-art VLM acceleration methods. Code will be made publicly available.
1 Introduction
(a) TwigVLM architecture.
(b) Two key advantages of TwigVLM.
Figure 1: Our TwigVLM overview. (a) Given a base deep VLM, the TwigVLM is obtained by freezing the whole VLM while training a shallow twig block upon an early layer of the VLM. (b) Compared to the VLM acceleration methods based on visual token pruning (e.g., FastV [8]), TwigVLM not only achieves better accuracy retention by employing a twig-guided token pruning (TTP) strategy, but also yields higher generation speed by utilizing a self-speculative decoding (SSD) strategy [12]. Both FastV and TwigVLM take LLaVA-1.5-7B [30] as the base VLM (red dotted lines in the bar graphs) with the same pruning ratio of 88.9%.
The revolution in large language models (LLMs) has fundamentally reshaped the landscape of artificial intelligence [4, 44, 17]. Meanwhile, there has been growing interest in building large vision-language models (VLMs) [39, 68, 29, 1, 9, 53, 25], which can perform general-purpose visual understanding and reasoning and be used in a wide range of tasks such as object grounding [40, 36], document understanding [54, 35], GUI controlling [47, 60], and robotics [3, 22]. Despite the remarkable progress achieved by these large VLMs, they are usually parametric-intensive and computational-heavy. This poses great challenges for using these models in computation-constrained platforms or low-latency scenarios. Therefore, the research on VLM acceleration is of practical and emergent demands.
As the computational cost of the transformer architecture [46] increase quadratically with the length of token sequence, an effective way to accelerate VLMs is to reduce the length of tokens. Moreover, visual tokens usually contain much more redundant information than textual tokens [8], especially in high-resolution images and long videos. Therefore, a series of works on visual token pruning has been proposed [8, 62, 32, 52]. For instance, FastV [8] leverages attention signals from an early layer of the VLM to drop redundant tokens for subsequent layers. Many following works have extended FastV’s idea to progressive token reduction across layers via pruning and merging operations [62, 32], while some others have been proposed to use the attention maps from the visual encoder to guide the visual token pruning before being input into the VLM backbone [52, 61]. Despite the success they have achieved, they still suffer from the following two major shortcomings:
  1. (i)
    To effectively reduce computational costs, a large proportion of visual tokens are usually pruned in the early layers, which is guided by the attention map of the same layer. However, the attention signals in the early layers are insensitive to the task, which restricts the quality of retained visual tokens and leads to a considerable accuracy drop of the accelerated VLM.
  2. (ii)
    The VLM inference process consists of the prefilling and decoding stages. Most token pruning-based approaches only focus on the acceleration of the prefilling stage while paying less attention to the decoding stage, which is the most time-consuming part of the inference process. This results in a significant efficiency gap between the theoretical estimation and practical application, especially in the tasks that generate long responses (e.g., more than 30 tokens).
We ask: Is it possible to address the two limitations above in one unified framework?
To this end, we present TwigVLM, a simple and general VLM acceleration approach by appending a lightweight twig block upon an early layer of the base VLM (see Figure 1(a)). After efficient post-training of the twig block only, the learned twig block is used in both the prefilling and decoding stages of inference. As shown in Figure 1(b), our TwigVLM not only achieves better accuracy retention by employing a twig-guided token pruning (TTP) strategy, but also yields higher generation speed by utilizing a self-speculative decoding (SSD) strategy [12]. To summarize, the synergistic effect of the TTP and SSD strategies enables holistic inference acceleration of both the prefilling and decoding stages at maximum accuracy retention.
Extensive experiments on a wide range of VLM benchmarks validate the effectiveness of the proposed method. Without bells and whistles, TwigVLM preserves 96% of the original accuracy after pruning 88.9% of visual tokens and achieves 154% speedup in generating long responses. These results demonstrate significant improvements over previous state-of-the-art methods in terms of both accuracy and speed. Moreover, our study also discloses a new aspect for VLM acceleration, i.e., long response generation, which is crucial in real-world scenarios.
2 Pilot Studies
Before introducing our TwigVLM, we conduct two pilot studies on the accuracy retention and generation speed, respectively, which reveals some crucial but easily overlooked observations in previous VLM acceleration works.
2.1 Preliminaries
To better understand the pilot studies, we first provide a concise overview of the architecture and two-stage inference process of modern VLMs. After that, we briefly introduce FastV [8], a typical VLM acceleration method via adaptive visual token pruning at inference.
VLM architecture and inference. Modern VLMs are usually comprised of a vision encoder and an LLM, where the LLM further consists of a stack of $L$ transformer layers.
During inference, the learned VLM first encodes an input image as a sequence of $M$ visual tokens $\mathbf{X_{v}}=[\mathbf{v}_{1},...,\mathbf{v}_{M}]\in\mathbb{R}^{M\times d}$ and represents a language prompt as a sequence of $N$ textual tokens $\mathbf{X_{q}}=[\mathbf{q}_{1},...,\mathbf{q}_{N}]\in\mathbb{R}^{N\times d}$, where $d$ is the common dimensionality of the multimodal tokens. These two groups of tokens are concatenated into $\mathbf{X}=[\mathbf{X_{v}},\mathbf{X_{q}}]$, which is then passed through the LLM to generate a sequence of $S$ response tokens $\mathbf{y}=(y_{1},...,y_{S})$ in an autoregressive manner. At each generation step $j$, we have:
$y_{j}=\mathop{\text{argmax}}\limits_{\hat{y}_{j}}p_{\mathrm{LLM}}(\hat{y}_{j}|\mathbf{X},\mathbf{y}_{<j}),$ (1)
where $\mathbf{y}_{<j}$ is the response sequence before the current token $y_{j}$. In practice, the iterative generation process is divided into the prefilling and decoding stages by leveraging the KV-cache technique [67].
Prefilling stage. This stage executes the forward pass of the input tokens $\mathbf{X}$, which remains unchanged during the response generation. Therefore, their intermediate representations in the LLM, namely the key-value pairs in each self-attention (SA) block of the transformer layer, can be cached and reused in the subsequent decoding stage. Specifically, the SA operation is given by:
$\mathrm{SA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathbf{A}\mathbf{V}=\mathrm{softmax}(\mathcal{M}(\mathbf{Q}\mathbf{K}^{T}/\sqrt{d}))\mathbf{V}$ (2)
$\mathbf{Q}=\mathbf{X}\mathbf{W_{q}},\quad\mathbf{K}=\mathbf{X}\mathbf{W_{k}},\quad\mathbf{V}=\mathbf{X}\mathbf{W_{v}},$
where $\mathcal{M}(\cdot)$ is the causal masking function and $\mathbf{A}$ is the attention map of query-key pairs. $\mathbf{Q},\mathbf{K},\mathbf{V}$ denote the queries, keys, and values, while $\mathbf{W_{q}},\mathbf{W_{k}},\mathbf{W_{v}}$ are their linear projection matrices, respectively.
As mentioned, the key-value pairs $\mathbf{K}$ and $\mathbf{V}$ of all transformer layers in the LLM are stored in the KV-cache to expedite the decoding stage.
Decoding stage. The generation of the response token $y_{j}$ is based on its previous tokens $(\mathbf{X},\mathbf{y}_{<j})$. As the attention computation in Eq.(2) is causal and parallel for each query, the generation for $y_{j}$ can be accelerated by feeding only the single token into the LLM and computing its corresponding key and value, while reusing the rest key-value pairs from the KV-cache. After each decoding step, the computed key-value pairs for $y_{j}$ are appended to the KV-cache for the efficient generation of subsequent tokens.
Figure 2: Attention quality for token selection. The RelAcc (defined in §4.2) is evaluated on GQA [19] and TextVQA [42], which measures the accuracy gap between LLaVA-1.5-7B and its FastV variants with $R$=128 and varied attention layer depth $D$. Taking an image with two prompts as an example, we visualize their attention maps from typical layers $D$=2 and 18, respectively.
Visual token pruning. The visual tokens in VLMs are much more than the textual tokens and exhibit severe redundancy [62, 32]. To accelerate the inference process of VLMs, some recent works such as FastV [8] and SparseVLM [62] have been proposed to drop redundant visual tokens based on the attention scores during the prefilling stage. Specifically, FastV works by passing the input tokens $\mathbf{X}$ through the first $K$ VLM layers to obtain the attention map $\mathbf{A}^{(K)}$ and token sequence $\mathbf{X}^{(K)}=[\mathbf{X}_{\mathbf{v}}^{(K)},\mathbf{X}_{\mathbf{q}}^{(K)}]$ after layer $K$. Denoting $R$ as the number of visual tokens after pruning, the retained token sequence $\mathbf{\hat{X}}^{(K)}$ is obtained as follows:
$\mathbf{\hat{X}}^{(K)}$ $=\mathcal{P}(\mathbf{X}^{(K)},\mathbf{A}^{K},R)$ (3)
$=\left[\text{TopR}\left(\mathbf{X}_{\mathbf{v}}^{(K)},\sum_{i=M+1}^{M+N}\mathbf{A}^{(K)}[i,1\text{:}M]\right),\mathbf{X}_{\mathbf{q}}^{(K)}\right].$
$\text{TopR}(\mathbf{X}^{(K)}_{\mathbf{v}},\mathbf{a})$ selects top-$R$ most significant visual tokens from $\mathbf{X}^{(K)}_{\mathbf{v}}$, which is guided by the attention scores $\mathbf{a}$ from textual tokens to visual tokens. In FastV, both $K$ and $R$ are set to small values, which means the token lengths are short in most VLM layers, thus the computational complexity of the prefilling stage is significantly reduced at the expense of a certain amount of performance drop. For a fair comparison of different token pruning methods, the average number of retained tokens $\bar{R}$ is used and defined as follows [^1] The computed result will be rounded to the nearest integer.:
$\bar{R}=[M\times K+R\times(L-K)]/L$ (4)
2.2 Study 1: Attention Quality for Token Selection
(a) Decoding time with varied $S$.
(b) P&D time at $S$=32
Figure 3: Prefilling and decoding time costs. (a) Prefilling time (gray dotted line) and decoding time for LLaVA-1.5-7B [30] with different response lengths. (b) Prefilling (P) and decoding (D) time comparisons of LLaVA-1.5-7B and its FastV-based variant.
The performance of visual token pruning methods heavily depends on the quality of the chosen attention map for pruning. Existing approaches often utilize the attention map from an early layer, e.g., $K$=2, to guide the visual token pruning at the same layer. We ask: Whether an attention map from a late layer contains more informative signals to facilitate token pruning? Notably, directly applying the attention map of a late layer to guide the token pruning at an early layer will inevitably introduce redundant computation, this pilot study only aims to evaluate the quality of the attention maps in different layers.
(a) Training stage
(b) Inference stage (prefilling)
(c) Inference stage (decoding)
Figure 4: Training and two-stage inference of TwigVLM. (a) The twig block is initialized from the base VLM and is coupled with the first $K$ layers of base VLM to form a shallow model, which can be trained efficiently. (b) In the prefilling stage, different from previous approaches that perform token selection based on the attention maps from the base VLM, TwigVLM introduces a twig-guided token pruning (TTP) strategy to obtain more precise signals to guide the pruning of visual tokens. (c) In the decoding stage, the shallow and deep sub-networks act as the draft and target models respectively, enabling it to perform the self-speculative decoding (SSD) [12] to accelerate the generation of long responses.
The pilot experiment is conducted as follows. We use LLaVA-1.5-7B as the base VLM and FastV as the token pruning strategy. Different from the original FastV that selects and prunes visual tokens at the same layer $K$, we perform token selection guided by the attention of the $D$-th layer and then perform token pruning at the $K$-th layer. In this experiment, we set $K$=2 and let $D\geq K$. From the results in Figure 2, we can see that the accuracy grows as the attention depth $D$ increases, showing that the attention maps from later layers provide more precise signals for visual token pruning than those from early layers. We hypothesize that later layers are closer to the prediction head (i.e., the loss function), thus their attention maps understand the relation between multimodal tokens more accurately. Moreover, from the visualized attention maps, we find that the selected tokens by early-layer attention ($D$=2) are insensitive to different prompts, while the selected tokens by late-layer attention ($D$=18) are more related to the prompt. This observation also verifies our hypothesis above.
2.3 Study 2: Prefilling and Decoding Time Costs
As mentioned above, the inference process consists of the prefilling and decoding stages. FastV and other token-pruning methods primarily accelerate the prefilling stage. We ask: Can these methods also facilitate the acceleration of the decoding stage, especially when the generated response is relatively long? This is a practical question since VLMs are often expected to generate long responses (e.g., 30 tokens) in real-world scenarios.
To answer this question, we conduct the pilot experiment as follows. We take LLaVA-1.5-7B [30] as the base VLM and compute its time spent on the prefilling and decoding stages for the response length $S$ ranging from 2 to 128. The results in Figure 3(a) show a linear increase in decoding time w.r.t. the response length $S$, while the prefilling time can be neglected when $S\geq$32.
Next, we compare the prefilling and decoding times of LLaVA-1.5-7B and its FastV-based variant (LLaVA+FastV) when generating responses of the same length $S$=32. From Figure 3(b), we observe that although FastV effectively reduce the prefilling time, it attains limited speedup in the decoding stage. This frustrating observation can be explained by two facts: (i) the KV-cache mechanism undermines the speedup for SA blocks achieved by FastV, and (ii) FastV does not accelerate the FFN blocks during decoding, but they account for the majority of the computational costs.
To summarize, these observations highlight the significance and necessity of improving the decoding efficiency for long response generation.
3 The Proposed TwigVLM
Inspired by the pilot studies above, we propose TwigVLM, a simple yet effective approach for accelerating VLMs. As depicted in Figure 4, our TwigVLM is obtained by training a lightweight twig block upon an early layer of a frozen VLM. Subsequently, the learned twig block is used in both the prefilling and decoding stages to simultaneously achieve more precise visual token pruning and faster generation speed for long responses.
3.1 Architecture and Training of TwigVLM
Denoting a $L$-layer base VLM as $\mathcal{M}_{b}=\{\mathcal{T}_{l}\}_{l=1}^{L}$, we attach a twig block $\{\mathcal{G}_{t}\}_{t=1}^{T}$, i.e., an extra lightweight module with $T$ transformer layer, upon an early layer of $\mathcal{M}_{b}$. Specifically, the twig block is appended after the $K$-th layer $\mathcal{T}_{K}$, resembling a twig growing from a tree trunk. When $K+T\ll L$, we obtain a very shallow VLM by combining the twig block and its based $K$ trunk layers $\mathcal{M}_{s}=\{\mathcal{T}_{k}\}_{k=1}^{K}\cup\{\mathcal{G}_{t}\}_{t=1}^{T}$. [^2] For notational simplicity, we omit the visual encoder and the prediction head in the definition of $\mathcal{M}_{b}$ and $\mathcal{M}_{s}$.
As shown in Figure 4(a), the training strategy of TwigVLM is simple yet efficient. All the model weights in the twig block are initialized from the layers $\{\mathcal{T}_{i}\}_{i=K+1}^{K+T}$ and the prediction head of $\mathcal{M}_{b}$, respectively. After that, the shallow VLM $\mathcal{M}_{s}$ is finetuned using the same training data and the conventional autoregressive (AR) loss as those for $\mathcal{M}_{b}$. Notably, only the lightweight twig block is updated during the TwigVLM training, consuming only about 10% of the training time compared to its base VLM.
3.2 Efficient inference with TwigVLM
The obtained TwigVLM can be seen as a composition of a deep VLM $\mathcal{M}_{b}$ and a shallow VLM $\mathcal{M}_{s}$, where their first $K$ layers are shared with each other. We can leverage this ingenious architecture to accelerate both the prefilling and decoding stages during inference.
Prefilling acceleration via token pruning. The prefilling stage can be accelerated by applying the visual token pruning strategy. Previous methods rely on attention maps from early layers to select key visual tokens and then prune the rest. As discussed in §$\S$2.2, we have verified that the attention map of a late layer, which is closer to the prediction head, can provide more precise signals than that of an early layer. Inspired by this observation, we introduce a twig-guided token pruning (TTP) strategy which utilizes the attention map from the last twig layer to select key tokens.
As shown in Figure 4(b), given the input tokens $\mathbf{X}$, the TTP strategy first feeds them through the first $K$ layers of $\mathcal{M}_{b}$ to obtain the latent representations $\mathbf{X}^{(K)}_{\mathcal{M}_{b}}$, and then feeds them through the twig block to obtain the attention map of the last twig layer $\mathbf{A}^{(K+T)}_{\mathcal{M}_{s}}$. After that, we use the token pruning function in Eq.(3) to obtain the representations of the retained visual tokens $\mathbf{\hat{X}}^{(K)}_{\mathcal{M}_{b}}$ as follows:
$\mathbf{\hat{X}}^{(K)}_{\mathcal{M}_{b}}=\mathcal{P}(\mathbf{X}^{(K)}_{\mathcal{M}_{b}},\mathbf{A}^{(K+T)}_{\mathcal{M}_{s}},R),$ (5)
where this condensed token sequence $\mathbf{\hat{X}}^{(K)}_{\mathcal{M}_{b}}$ is further fed through the remaining layers of $\mathcal{M}_{b}$.
To achieve aggressive token pruning with small $\bar{R}$, both $K$ and $R$ need to be small values according to Eq.(4), leading to considerable performance drops. To mitigate this issue, several works adopt progressive token pruning strategies [62, 50] to retain more tokens at middle layers. In addition, previous studies have shown that the visual tokens in late layers (e.g., after the 20th layer) barely contribute to the prediction [65, 64]. Motivated by these works, we additionally introduce a simple FinalWipe strategy by removing all the visual tokens after the $K_{\text{f}}$-th layer (a late layer) in $\mathcal{M}_{b}$. By doing so, Eq.(4) can be reformulated as follows:
$\bar{R}=[M\times K+R\times(K_{\text{f}}-K)]/L.$ (6)
When $\bar{R}$ and $K$ are fixed, this strategy enables a larger $R$, which facilitates the performance of our TwigVLM.
Decoding acceleration via self-speculation. The acceleration for the decoding stage remains underexplored, especially for the long response scenarios. TwigVLM’s architecture, which contains two sub-networks, namely the deep one $\mathcal{M}_{b}$ and shallow one $\mathcal{M}_{s}$, enables the self-speculative decoding (SSD) strategy [12, 28] to address this challenge. As illustrated in Figure 4(c), we take the shallow model $\mathcal{M}_{s}$ as a draft model to efficiently generate multiple subsequent tokens (i.e., draft tokens), and then take the deep model $\mathcal{M}_{b}$ as a target model to verify these tokens in parallel.
Specifically, SSD operates through multiple draft-then-verify iterations to generate the response. In each iteration, the draft model generates multiple tokens in an autoregressive manner. These tokens are then verified by feeding them through the target model in a single forward pass to determine their acceptance. Although the total computational costs are not reduced, the SSD strategy accelerates the decoding stage by processing multiple draft tokens in parallel, which maximally utilizes GPUs’ parallel computing capabilities. Moreover, the draft and target models in TwigVLM share the computation and KV-cache of the first $K$ layers, thus achieving further efficiency gains.
It is worth noting that the SSD strategy produces exactly the same response as the target model, which means the accuracy of TwigVLM is only affected by the TTP strategy.
4 Experiments
We evaluate the performance of TwigVLM on three popular VLMs, namely LLaVA-1.5-7B [30], LLaVA-NeXT-7B [31] and Video-LLaVA-7B [26], to compare with the state-of-the-art VLM acceleration methods. After that, we conduct comprehensive ablation studies to analyze the effectiveness of the key elements of TwigVLM.
4.1 Implementation Details
Unless otherwise noted, we implement TwigVLM with the following hyper-parameters: the number of twig layers $T$=3, the token pruning position $K$=2, and the FinalWipe position $K_{\text{f}}$=24. Based on the default setting above, we adjust $R$ using Eq.(6) to fairly compare TwigVLM with other methods at the same pruning ratio $1-\bar{R}/M$. All experiments are conducted on a server with 8$\times$×A100 GPUs. More implementation details are provided in the supplementary.
4.2 Evaluation Metrics
To evaluate the performance of TwigVLM, we consider two key aspects: accuracy preservation and inference speedup, which are respectively measured by the two metrics as follows. Relative Accuracy (RelAcc) is a widely used metric in previous works [8, 52], which calculates the proportion of the accuracy of the pruned model to the base model. When multiple benchmarks are provided, their relative accuracies are separately calculated and then averaged as the final RelAcc. Relative Speed (RelSpd) quantifies the relative generation speed of a pruned model compared to its base model. The generation speed is computed by dividing the number of generated tokens by the generation time—which encompasses both the prefilling and decoding stages—and then averaged across samples from a specified benchmark.
[…]
Upper Bound, 576 Tokens  (100%)
  LLaVA-1.5-7B 61.9 64.7 1862 58.2 69.5 78.5 100%
[…]
FastV [8] 56.5 63.7 1786 57.3 69.5 74.6 96.5%
SparseVLM [62] 57.6 62.5 1721 56.1 69.1 75.6 95.7%
PDrop [50] 57.3 63.3 1797 56.5 69.2 75.1 96.5%
MustDrop [32] 58.2 62.3 1787 56.5 69.2 76.0 96.6%
VisionZip [52] 59.3 63.0 1783 57.3 68.9 76.8 97.4%
VisionZip‡ [52] 60.1 63.4 1834 57.8 68.2 77.4 98.3%
TwigVLM (ours) 61.2 64.0 1848 58.0 68.8 78.1 99.2%
[…]
FastV 53.0 61.4 1646 56.0 69.5 69.2 92.2%
SparseVLM 56.0 60.0 1696 54.9 67.1 73.8 93.2%
PDrop 57.1 61.6 1761 56.6 68.4 72.9 95.1%
MustDrop 56.9 61.1 1745 56.3 68.5 74.6 95.1%
VisionZip 57.6 62.0 1762 56.8 68.9 75.6 96.1%
VisionZip‡ 58.9 62.6 1823 57.0 68.3 76.6 97.3%
TwigVLM (ours) 60.6 63.5 1818 57.8 69.5 77.9 98.7%
[…]
FastV 44.1 45.9 1218 50.7 70.0 52.0 77.0%
SparseVLM 52.7 56.2 1505 51.8 62.2 68.2 89.9%
PDrop 47.5 58.8 1561 50.6 69.0 69.2 87.6%
FasterVLM [61] 51.5 58.5 1573 53.1 69.6 66.8 89.1%
MustDrop 53.1 60.0 1612 54.2 63.4 69.3 89.6%
VisionZip 55.1 60.1 1690 55.5 69.0 72.4 93.3%
VisionZip‡ 57.0 61.5 1756 56.0 68.8 74.2 95.2%
TwigVLM (ours) 58.8 60.4 1760 55.8 70.0 75.6 96.0%
Table 1: Performance comparisons with three pruning ratios on six VLM benchmarks [19, 33, 13, 42, 34, 15]. All methods are applied on the same base model LLaVA-1.5-7B. The best result for each benchmark and pruning ratio is bolded.
[…]
Upper Bound, 2880 Tokens  (100%)
LLaVA-NeXT-7B 64.2 67.9 1851 61.3 70.2 81.8 100%
[…]
FastV 62 65.8 1807 60 69.1 79.5 97.6%
SparseVLM 60.3 65.7 1772 57.8 67.7 77.1 95.4%
VisionZip 61.3 66.3 1787 60.2 68.1 79.1 97.1%
VisionZip‡ 62.4 65.9 1778 60.8 67.9 79.9 97.4%
TwigVLM (ours) 63.4 67.4 1864 58.6 69.9 81.2 99.0%
[…]
FastV 54.9 60 1539 54.8 68.2 69.6 88.2%
SparseVLM 57.7 64.3 1694 55.9 67.3 73.4 92.3%
MustDrop 57.3 62.8 1641 59.9 68.0 73.7 92.6%
VisionZip 59.3 63.1 1702 58.9 67.3 76.2 93.8%
VisionZip‡ 61.0 64.4 1770 59.3 67.5 78.4 95.8%
TwigVLM (ours) 62.2 65.0 1758 57.4 68.7 79.7 96.2%
Table 2: Performance comparisons on LLaVA-NeXT-7B. This table follows the same layout and notes as Table 1.
Method TGIF     MSVD MSRVTT ActivityNet RelAcc
Video-LLaVA-7B 47.1 69.8 56.7 43.1 100.0%
FastV 23.1 38.0 19.3 30.6 52.1%
SparseVLM 44.7 68.2 31.0 42.6 86.5%
VisionZip 42.4 63.5 52.1 43.0 93.2%
TwigVLM(ours) 44.7 68.3 54.6 41.5 96.3%
Table 3: Performance comparisons on four widely used VideoQA benchmarks [20, 51, 58]. The base model is Video-LLaVA-7B, which originally has 2,048 visual tokens. We follow the compared methods to retain 135 visual tokens.
4.3 Main Results
Results on LLaVA-1.5 and LLaVA-NeXT. As shown in Tables 1 and 2, we take LLaVA-1.5-7B and LLaVA-NeXT-7B as the base VLMs and compare TwigVLM with the state-of-the-art VLM acceleration methods on six commonly-used VLM benchmarks. We can see that TwigVLM consistently and significantly outperforms all its counterparts under different pruning ratios. Notably, TwigVLM achieves near-perfect performance preservation (more than 99.0%) when pruning 66.7% visual tokens on LLaVA-1.5-7B and 77.8% tokens on LLaVA-NeXT-7B, showing the superiority of our TTP strategy over existing pruning strategies. Comparisons on more benchmarks and base VLMs are provided in the supplementary.
To better understand the effectiveness of our approach, we compare TwigVLM with two representative methods [8, 52] by visualizing the attention map used for token selection. The examples in Figure 5 suggest that TwigVLM can better understand the fine-grained semantics in both the prompt and image, thus identifying more informative visual tokens for token pruning.
[…]
Figure 5: Visualized attention map comparisons of TwigVLM and two typical token pruning methods. The visualized attention maps show that TwigVLM identifies accurate visual tokens to the prompt and predicts the right answer, while both counterparts fail to do that. More examples are provided in the supplementary.
attention source, depth $D$ RelAcc (%)
(a) VLM backbone, $K$ 82.3
(b) VLM backbone, $K$+$T$ 86.2
(c) twig layer, $K$+$T$ 96.0
(a) Visual token selection. The attention map from the last twig layer results in the highest accuracy as it is close to the prediction head.
acceleration strategy RelSpd (%)
(a) token purning (FastV [8]) 104.3
(b) SSD 146.7
(c) TTP & SSD 153.6
(b) Acceleration strategies. The TTP and SSD strategies are complementary and their synergy achieves the highest speedup.
initialization strategy RelAcc (%) RelSpd (%)
(a) random init. 87.2 120.4
(b) VLM layers[$L$-$T$:$L$] 90.4 131.4
(c) VLM layers[$K$:$K$+$T$] 96.0 153.6
(c) Twig block initialization. The twig layers initialized from the $K$-th to ($K$+$T$)-th layers in base VLM achieves the best accuracy and speed.
     $T$   RelAcc (%)   RelSpd (%)
1 93.9 154.1
2 95.2 152.6
3 96.0 153.6
4 95.8 145.4
(d) Number of twig layers. $T$=3 results in the highest accuracy with negligible speed degradation compared to $T$=1.
  $K$   $R$ RelAcc (%) RelSpd (%)
0 85 93.6 137.5
1 64 95.5 136.5
2 41 96.0 153.6
3 15 85.5 145.2
(e) Pruning position. When $\bar{R}$ is fixed to 64, $K$=2 achieves the optimal balance between accuracy and speed.
FinalWipe $K_{\text{f}}$ $R$ RelAcc (%) RelSpd (%)
$\times$ 32 30 93.1 154.6
20 50 95.8 151.3
24 41 96.0 153.6
28 37 95.1 154.1
(f) FinalWipe. The FinalWipe strategy facilitates model accuracy and $K_{\text{f}}$=24 is the optimal choice to balance accuracy and speed.
Table 4: Ablation experiments for TwigVLM. Taking LLaVA-1.5-7B as the base VLM, the relative accuracy (RelAcc) is evaluated on the six benchmarks mentioned in Table 1 and the relative speed (RelSpd) is evaluated on MM-Vet. In each table, the best result is bolded and the result with the default setting ($T$=3, $K$=2, $K_{\text{f}}$=24, $\bar{R}$=64) is marked in green.
Results on Video-LLaVA for video understanding. Our TwigVLM is general and can be extended to the challenging video domain seamlessly. As shown in Table 3, TwigVLM consistently surpasses all the state-of-the-art counterparts on most VideoQA benchmarks and achieves 3.1% improvements over VisionZip in terms of RelAcc. These results further validate the generalization ability of our TwigVLM.
Figure 6: Generation speed comparisons on two benchmarks: (a) TextVQA with short responses and (b) MM-Vet with long responses. $\bar{S}$ denotes the average number of generated tokens on the whole benchmark. The RelSpd of each bar is highlighted in red.
Generation speed comparisons. As mentioned above, TwigVLM can effectively accelerate the generation. To validate this, we conduct intensive experiments on two typical benchmarks TextVQA [42] and MM-Vet [57] to compare the speedup of different VLM acceleration methods based on LLaVA-1.5-7B. The results in Figure 6 show that: (i) TwigVLM achieves superior or competitive speedup in all configurations, suggesting dual advantages in speed and accuracy. (ii) All the methods attain the similar RelSpd (120$\sim$130%) on TextVQA with short responses. In contrast, TwigVLM delivers significantly higher speedup ($\sim$150%) than FastV ($\sim$104%) and VisionZip ($\sim$106%) on MM-Vet with long responses. This reveals the superiority of our SSD strategy in long response generation. (iii) A higher pruning ratio (i.e., a smaller $\bar{R}$) leads to more speedup on TextVQA, but has little effect on MM-Vet, which has been observed and explained in §$\S$2.3.
4.4 Ablation Studies
To validate the effectiveness of TwigVLM’s key components, we conduct ablation experiments using the default setting ($T$=3, $K$=2, $K_{\text{f}}$=24, $\bar{R}$=64). Taking LLaVA-1.5-7B as the reference model, the relative accuracy (RelAcc) is evaluated on the six benchmarks mentioned in Table 1 and the relative speed (RelSpd) is evaluated on MM-Vet. Results in Table 4 are discussed in detail below.
Visual token selection. The performance of visual token pruning relies on the chosen attention map for token selection. Table 4(a) compares three TwigVLM variants with token pruning guided by different attention sources: (a) the $K$-th layer of the VLM backbone, (b) the ($K$+$T$)-th layer of the VLM backbone, and (c) the last twig layer which has the same depth as (b). From the results, we can see that using the attention map of the $K$-th backbone layer yields the lowest accuracy (82.3%), which verifies the conclusion in §$\S$2.3 that the semantic in early-layer attention is ambiguous. With the same attention depth $D$=($K$+$T$), selecting tokens by the attention from the backbone (86.2%) is inferior to that from the twig layer (96.0%), indicating that the quality of the attention map is also impacted by the “distance” between the layer it is on and the prediction head.
Acceleration strategies. The inference acceleration of TwigVLM is achieved by the synergy between the twig-guided token pruning (TTP) and self-speculative decoding (SSD) strategies during prefilling and decoding, respectively. In Table 4(b), we compare the three acceleration strategies as follows: (a) visual token pruning by FastV [8], (b) SSD only without token pruning, and (c) the standard TwigVLM with SSD and TTP. While FastV achieves modest speedup (104.3%), SDD delivers a remarkable improvement (146.7%) as the decoding time is longer than prefilling. Finally, the TTP and SSD strategies are complementary and their synergy achieves the highest speedup (153.6%).
Twig block initialization. The initialization strategy for twig layers has a prominent impact on TwigVLM’s performance. From Table 4(c), we can see that: random initialization in (a) performs poorly due to the lack of knowledge inherited from the base VLM. Moreover, the strategy in (b) that uses the last $T$ VLM layers to initialize the twig layers results in distinct improvements. Finally, initializing from layers $K$ to $K+T$ achieves the best results in both accuracy and speed. The speed improvements are contributed by the higher token acceptance rate [^3] The token acceptance rate measures how often verification accepts each of the draft tokens [12].. More detailed analyses about token acceptance rate are provided in the supplementary.
Number of twig layers. Table 4(d) varies the number of twig layers $T$ from 1 to 4 to investigate its effect on TwigVLM. We observe that a single twig layer ($T$=1) leads to the highest speed but the lowest accuracy. Increasing $T$ to 2 or 3 improves accuracy with negligible speed degradation. However, further increasing $T$ to 4 leads to a distinct performance drop in terms of both accuracy and speed. This suggests that $T$=3 achieves an optimal balance in providing more accurate attention signals for pruning (via more twig layers) while still maintaining computational efficiency due to the high token acceptance ratio in SSD.
Pruning position. To achieve an optimal number of retained tokens when $\bar{R}$ is fixed to 64, we use Eq.(6) to explore different combinations of $K$ and $R$. Increasing $K$ leads to a smaller $R$, which means a trade-off between the number of visual tokens in early layers and the number of visual tokens in subsequent layers. As shown in Table 4(e), both the accuracy and speed achieve the best results at $K$=2, which is chosen in our default setting.
FinalWipe. The FinalWipe works by removing all visual tokens after the $K_{\text{f}}$-th layer. Table 4(f) ablates different TwigVLM variants without/with this strategy. When FinalWipe is not introduced, the accuracy is relatively low due to the insufficiency of retained tokens. Introducing FinalWipe facilitates accuracy with limited speed drop, which can be explained by the fact that the visual tokens become less important in late layers. $K_{\text{f}}$=24 achieves a good balance between accuracy and speed, thus is used as default.
5 Related Work
Inference acceleration for LLMs. Accelerating the inference process of LLMs has attracted significant attention from both academia and industry. Although various efficient model architectures have been proposed successively [16, 43, 27], they often need to be trained from scratch, thus incurring substantial costs. Consequently, the research of accelerating off-the-shelf LLMs has been extensively investigated from the prefilling and decoding stages of inference, respectively. In the prefilling stage, acceleration is achieved by redundant token reduction [21, 38] and attention operator optimization [11, 10]. In the decoding stage, techniques such as dynamic reduction of KV-cache [49, 63] and speculative decoding [23, 5, 28, 12] are introduced. In particular, speculative decoding performs parallel verification of draft tokens to boost GPU utilization, thus significantly improving the speed.
Visual token reduction for VLMs. Unlike acceleration techniques for LLMs, token reduction methods for VLMs mainly address the problem of visual token redundancy to improve inference efficiency [65]. The pioneering work of FastV [8] first studies this redundancy phenomenon, revealing that pruning visual tokens at an early layer of the VLM can achieve acceleration while preserving the majority of the model’s accuracy. Following this paradigm, subsequent studies have introduced adaptive or hierarchical pruning strategies [62, 55, 32, 56, 48, 18] and training-based adoption [7, 50, 64] to further reduce the accuracy drop. Meanwhile, there is another line of studies that performs more aggressive token reduction by pruning or merging visual tokens in the visual encoding stage [2, 41, 52, 61, 45]. However, the accuracy of these pruned models is suboptimal as they are unaware of the textual prompt when pruning visual tokens. Finally, although the above methods effectively accelerate the prefilling stage of VLMs, their speedup in the decoding stage is limited.
Collaborative decoding of small-large VLMs. Beyond visual token reduction, some recent works have explored another way, i.e., the collaborative decoding of small and large VLMs, for VLM acceleration [14, 66]. Specifically, Gagrani et al. introduce a multimodal speculative decoding method, where a small language-only LLM is adopted to efficiently generate draft tokens and followed by the verification by a large VLM [14]. Similarly, Zhao et al. use an small VLM for early-exit prediction, invoking the large VLM only when the prediction confidence is low [66]. However, both methods above require an pre-existing and high-capable small model with the same vocabulary as the large model, which severely restricts the choice of VLMs.
6 Conclusion
We present TwigVLM, a conceptually simple yet effective approach for VLM acceleration. Unlike existing purely token pruning methods, TwigVLM trains a lightweight twig block upon an early layer of the base VLM. After that, the learned twig block is used in both the prefilling and decoding stages to obtain better accuracy retention and faster generation speed simultaneously. Extensive ablation, comparative experiments, and comprehensive analyses on a wide range of VLM benchmarks show the superiority of TwigVLM over previous state-of-the-art methods. We hope our work may serve as a solid baseline to inspire future research on VLM acceleration.
Appendix A More Implementation Details
TwigVLM training. As described in the main text, the twig block is trained by finetuning the shallow VLM $\mathcal{M}_{s}$. Specifically, $\mathcal{M}_{s}$ is initialized with the weights of the first $K$+$T$ layers and the prediction head of the corresponding base VLM $\mathcal{M}_{b}$. During finetuning, only the last $T$ layers and the prediction head—collectively termed the twig block—are updated, while the first $K$ layers remain frozen. This process follows the same training manner to train the base VLM $\mathcal{M}_{b}$. Theoretically, any suitable multimodal instruction tuning dataset can be employed to finetune $\mathcal{M}_{s}$.
In our experiments, we utilize the LLaVA-665K dataset [30] to train the twig blocks for the LLaVA-1.5-7B and LLaVA-NeXT-7B models, and combine the LLaVA-665K and VideoInstruct-100K datasets [26, 37] for the Video-LLaVA-7B model. The optimization hyper-parameters are detailed in Table 5. All training is performed on a server equipped with 8 NVIDIA A100 GPUs. Under these conditions, the training process is highly efficient, requiring only approximately 10% of the time needed to train the corresponding base VLM, e.g., training the twig block for the LLaVA-1.5-7B model takes about 1 hour, while the training of the original LLaVA-1.5-7B takes about 12 hours.
config setting
optimizer AdamW
weight decay 0.
optimizer momentum $\beta_{1},\beta_{2}{=}0.9,0.98$
batch size 128
learning rate schedule cosine decay
peak learning rate 5e-5
warm-up strategy linearly warm-up
warm-up ratio 0.03
training epochs 1
Table 5: Training settings. These hyper-parameters are shared across all the TwigVLM models in our experiments.
Twig-guided token pruning (TTP). During inference, TwigVLM leverages the TTP strategy to perform token pruning over the base VLM: (i) at the $K$-th layer, selecting $R$ key visual tokens (output by the $K$-th layer) and discarding the rest tokens guided by the attention map from the last twig layer, and (ii) applying the FinalWipe strategy to further remove all the visual tokens after the $K_{\text{f}}$-th layer. Therefore, we adjust the value of $R$ to satisfy different pruning ratios calculated by the average number of retained visual tokens $\bar{R}$. Table 6 shows the default pruning settings for TwigVLM under different pruning ratios.
$\bar{R}$ pruning ratio $K$ $R$ $K_{\text{f}}$
192 66.7% 2 227 24
128 77.8% 2 134 24
64 88.9% 2 41 24
Table 6: Pruning settings. These hyper-parameters correspond to the default TTP settings of different pruning ratios.
Self-speculative decoding (SSD). For efficient generation of long responses, TwigVLM applies the SSD strategy by using the $\mathcal{M}_{s}$ as the draft model and $\mathcal{M}_{b}$ as the target model. Specifically, in each SSD iteration, the draft model efficiently predicts $\delta=5$ subsequent draft tokens in an autoregressive manner. To further improve efficiency, this draft generation process is equipped with an early-exit mechanism that allows the draft model to stop generation if the probability of the current predicted token falls below a predefined threshold $\theta=0.6$. The target model then verifies these generated draft tokens in parallel, accepts those matching the target model’s predictions, and then predicts a next token by itself. The iteration repeats until the <EOS> token is generated. Note that the TTP and SSD strategies can be seamlessly integrated, as detailed in Algorithm 1.
Algorithm 1 Pseudocode of TwigVLM’s inference process
# bVLM: the base VLM model, i.e., M_b
# twig: the twig block
# K: Number of shared low layers
# K_f: The position to apply FinalWipe
# R: Number of retained visual tokens when pruning
# delta: Maximum draft token length
# theta: Confidence threshold to stop draft
def sVLM_forward(tokens):
X_k = bVLM.forward_low_layers(tokens, k=K)
prob, Attn_last = twig.forward(X_k)
a_i = argmax(prob)
return X_k, prob, Attn_last, a_i
def TwigVLM_inference(img, ques):
draft_toks = [] # temporary buffer for draft tokens
final_resp = [] # buffer for final response
# Prefilling stage of sVLM
X_k, _, Attn, a_i = sVLM_forward((img, ques))
draft_toks.append(a_i)
# Prune visual tokens in X_k using Eq. (5)
# X_k_b means shared token latents for bVLM
X_k_b = pruning(X_k, Attn, r=R)
# The loop of self speculative decoding
while EOS_TOKEN not in final_resp:
X_k, prob, _, a_i = sVLM_forward(a_i)
draft_toks.append(a_i)
X_k_b = concat(X_k_b, X_k, axis=1)
# the condition to stop draft and verify
if len(draft_toks) >= delta or prob < theta:
# removing all visual tokens after layer K_f
tgt_probs = bVLM.forward_high_layers(
X_k_b, k=K, fianl_wipe=K_f)
# verification
right_toks = [a for a, p in zip(draft_toks, tgt_probs[:-1]) if argmax(p) == a]
right_toks.append(argmax(tgt_probs[-1]))
final_resp.extend(right_tokens)
# reset temporary variables
draft_toks = []
X_k_b = None
a_i = final_resp[-1]
return final_resp
ablation variant   TokAR (%) RelSpd (%)
Twig block initialization (Table 4c in main text)
(a) random init. 37.7 120.4
(b) VLM layers[$L$-$T$:$L$] 44.1 131.4
(c) VLM layers[$K$:$K$+$T$] 57.4 153.6
Number of twig layers (Table 4d in main text)
(d) $T$ = 1 48.7 154.1
(e) $T$ = 2 53.4 152.6
(f) $T$ = 3 57.4 153.6
(g) $T$ = 4 58.1 145.4
Table 7: Token acceptance rate in SSD. We evaluate the token acceptance rate (TokAR) of the variants in the ablation experiments of the main text.
Appendix B More Experimental Results
Token acceptance rate in SSD. In the context of speculative decoding methods [23, 28, 12], the token acceptance rate (abbr. TokAR) serves as a critical metric for assessing the efficacy of these approaches. TokAR is defined as the proportion of the draft tokens generated by the draft model that are subsequently accepted by the target model. In TwigVLM, TokAR plays a key role, which is influenced by the effectiveness of the twig block and has a significant impact on model’s generation speed.
To analyze how TokAR is influenced by the design choices in TwigVLM, we evaluate this metric on several representative variants from the ablation studies presented in the main text. From the results shown in Table 7, we have the following findings: (i) A more effective draft model can be trained by only modifying the initialization strategy without altering the architecture. The variant (c) achieves the highest TokAR (57.4%) and thus the highest generation speedup. (ii) Increasing the number of twig layers $T$ introduces more computational costs while improving TokAR at the same time. As a result, the RelSpd exhibits only a modest decline when $T$ increases from 1 to 3. However, it drops distinctly at $T$=4, which indicates that TokAR begins to saturate. These findings suggest that TwigVLM achieves higher speedup by striking an optimal balance between TokAR and computation costs of the draft model.
[…]
Upper Bound, 576 Tokens  (100%)
LLaVA-1.5-13B 63.2 67.7 1818 61.3 72.8 80.0 100%
[…]
FastV 60.3 67.4 1807 60.4 74.0 77.7 98.6%
VisionZip 59.1 66.9 1754 59.5 73.5 78.1 97.4%
VisionZip‡ 61.6 67.1 1790 59.9 72.7 78.6 98.5%
TwigVLM (ours) 62.5 68.6 1840 60.4 73.1 79.4 99.9%
[…]
FastV 57.5 65.9 1758 58 73.8 74.3 95.7%
VisionZip 57.9 66.7 1743 58.7 74.0 76.8 96.6%
VisionZip‡ 60.1 67.6 1736 59.2 73.0 77.6 97.4%
TwigVLM (ours) 61.2 66.9 1811 60.2 73.4 79.1 98.9%
[…]
FastV 50.1 55.9 1408 52.2 73.2 61.1 83.6%
VisionZip 56.2 64.9 1676 57.4 74.4 73.7 94.2%
VisionZip‡ 58.1 65.6 1671 58.5 72.3 75.2 94.9%
TwigVLM (ours) 60.0 67.4 1765 58.4 72.4 77.0 97.1%
Table 8: Performance comparisons of TwigVLM with other token pruning methods on LLaVA-1.5-13B.
[…]
Upper Bound, 576 Tokens  (100%)
LLaVA-1.5-7B 61.9 64.7 1862 58.2 69.5 78.5 85.9 36.3 31.1 100%
[…]
FastV [8] 56.5 63.7 1786 57.3 69.5 74.6 79.2 35.7 28.1 95.6%
SparseVLM [62] 57.6 62.5 1721 56.1 69.1 75.6 83.6 33.8 31.5 96.2%
PDrop [50] 57.3 63.3 1797 56.5 69.2 75.1 82.3 - - 96.4%
MustDrop [32] 58.2 62.3 1787 56.5 69.2 76.0 82.6 - - 96.6%
VisionZip [52] 59.3 63.0 1783 57.3 68.9 76.8 85.3 36.6 31.7 98.5%
VisionZip‡ [52] 60.1 63.4 1834 57.8 68.2 77.4 84.9 36.2 32.6 99.2%
TwigVLM (ours) 61.2 64.0 1848 58.0 68.8 78.1 87.2 36.6 34.1 100.8%
[…]
FastV 53.0 61.4 1646 56.0 69.5 69.2 73.2 36.3 28.0 92.1%
SparseVLM 56.0 60.0 1696 54.9 67.1 73.8 80.5 33.8 30.0 93.6%
PDrop 57.1 61.6 1761 56.6 68.4 72.9 82.3 - - 95.2%
MustDrop 56.9 61.1 1745 56.3 68.5 74.6 78.7 - - 94.6%
VisionZip 57.6 62.0 1762 56.8 68.9 75.6 83.2 37.9 32.6 98.1%
VisionZip‡ 58.9 62.6 1823 57.0 68.3 76.6 83.7 37.3 32.9 98.8%
TwigVLM (ours) 60.6 63.5 1818 57.8 69.5 77.9 86.6 36.6 32.8 99.9%
[…]
FastV 44.1 45.9 1218 50.7 70.0 52.0 55.6 34.0 17.8 75.3%
SparseVLM 52.7 56.2 1505 51.8 62.2 68.2 75.1 32.7 23.3 85.6%
PDrop 47.5 58.8 1561 50.6 69.0 69.2 55.9 - - 84.4%
FasterVLM [61] 51.5 58.5 1573 53.1 69.6 66.8 67.2 - 27.5 87.6%
MustDrop 53.1 60.0 1612 54.2 63.4 69.3 68.0 - - 88.1%
VisionZip 55.1 60.1 1690 55.5 69.0 72.4 77.0 36.2 31.7 94.5%
VisionZip‡ 57.0 61.5 1756 56.0 68.8 74.2 80.9 35.6 30.2 95.6%
TwigVLM (ours) 58.8 60.4 1760 55.8 70.0 75.6 82.7 35.9 31.1 96.8%
Table 9: Performance of TwigVLM on LLaVA-1.5-7B compared to existing methods under three different pruning ratios. The best result for each benchmark and pruning ratio is bolded.
More accuracy comparisons on benchmarks. To further demonstrate the superiority of TwigVLM, we present additional experimental results on a larger base VLM and more benchmarks, respectively.
Table 8 provides the accuracy comparisons on a larger LLaVA-1.5-13B model. TwigVLM consistently achieves the best overall RelAcc compared to all the counterparts, with its superiority being more significant as the increase of pruning ratios. These results verify the scalability and generalization ability of TwigVLM in accelerating large VLMs.
Taking LLaVA-1.5-7B as the base VLM, Table 9 compares the accuracies between TwigVLM and other visual token pruning methods on nine VLM benchmarks under three different pruning ratios. TwigVLM consistently outperforms or matches its counterparts on all benchmarks and pruning ratios, achieving the best overall RelAcc. In particular, TwigVLM even surpasses the upper bound given by the base VLM in RelAcc (100.6%) with a 66.7% pruning ratio, demonstrating its spectacular effectiveness and robustness in accelerating VLMs to deal with various tasks.
[…]
Figure 7: Visualization of attention maps and predictions for FastV [8], VisionZip [52], and our TwigVLM on the examples chosen from the GQA [19] and TextVQA [42]. For the examples in the last row, we additionally provide the predictions from the LLaVA-1.5-7B.
Figure 8: Examples of the generated responses using the self-speculative decoding (SSD) on MM-Vet [57], with accepted tokens by the target model being highlighted in green.
Appendix C More Visualized Results
In this section, we provide more visualized results to validate the effectiveness of TwigVLM’s two key components: the twig-guided visual token pruning (TTP) and self-speculative decoding (SSD). We use LLaVA-1.5-7B as the base VLM in the following experiments.
Visual token pruning. To better understand the effectiveness of the proposed TTP strategy, we compare TwigVLM with two representative token pruning methods, namely FastV [8] and VisionZip [52], by visualizing their attention map for token selection and providing the corresponding answer predictions. We provide 16 examples from the GQA and TextVQA benchmarks. As illustrated in Figure 7, TwigVLM demonstrates superior ability to comprehend the semantics in both the textual prompt and image, and accurately identify task-specific image patches (i.e., visual tokens), thereby activating more informative visual tokens for token pruning. In contrast, FastV and VisionZip often fail to capture the fine-grained visual details, leading to suboptimal token selection and incorrect predictions. Notably, even though TwigVLM predicts an incorrect answer, its activated visual tokens according to the attention map is reasonable. This suggests that TwigVLM’s occasional failures may not be caused by the visual token pruning, but due to the limitations of the base VLM. These findings verify and explain the effectiveness of the TTP strategy.
Self-speculative decoding. To better understand the decoding behavior of the SSD strategy in TwigVLM, we show 8 examples of generated long responses on MM-Vet. From the results in Figure 8, we have two key observations: (i) In general, the proportion of accepted tokens (in green) surpasses that of the corrected tokens (in black) by the target model, indicating that TwigVLM achieves significant speedup through its high token acceptance rate. (ii) The majority of easy tokens, such as those associated with grammar and punctuation, are readily accepted. In contrast, the hard tokens, which often demand complex reasoning, have a high probability to be corrected by the target model. In practice, the proportion of easy tokens is usually larger than the hard ones, which confirms the the effectiveness of our SSD strategy in accelerating the decoding stage while maintaining the generation quality.
Appendix D Evaluation Benchmarks
In this section, we provide a brief overview of the benchmarks used in our experiments.
GQA [19] is a benchmark that focuses on visual scene understanding and reasoning, leveraging scene graphs, questions, and images. It incorporates spatial relationships and object properties, posing challenges for models to perform accurate visual reason under complex visual environments.
MMBench [33] adopts a hierarchical evaluation approach with three levels: Level-1 (perception and reasoning), Level-2 (six sub-abilities), and Level-3 (20 specific dimensions). This structured framework allows for a comprehensive evaluation of model performance, making it an effective tool for assessing a wide range of visual understanding capabilities. We denote it as “MMB” in the main text.
MME [13] assesses models across 14 subtasks that probe both perceptual and cognitive skills. Carefully crafted instruction-answer pairs guarantee a fair and comprehensive evaluation of a model’s multimodal performance. The final score reported on this benchmark is the summation of both the perception and cognition scores.
[…]
VQA-v2 [15] is a large-scale benchmark featuring 265K images of real-world scenes and objects, with each image paired with open-ended questions and 10 human-provided ground truth answers.
[…]
POPE [24] targets object hallucination evaluation by posing binary questions about object presence in images. It employs metrics such as Accuracy, Recall, Precision, and F1 score across three sampling methods. The reported score is calculated by the mean accuracy over the three indicators: adversarial, random, and popular.
MMMU [59] challenges models with tasks requiring college-level expertise and reasoning skills. It comprises 11.5K questions drawn from exams, quizzes, and textbooks, spanning six key disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering. Featuring 30 subjects and 183 sub-fields, MMMU involves diverse image types, e.g., charts, diagrams, and chemical structures, demanding advanced perceptual and domain-specific reasoning abilities akin to those of human experts.
MM-Vet [57] evaluates six fundamental vision-language capabilities: recognition, OCR, knowledge, language generation, spatial awareness, and mathematical reasoning. It examines 16 specific combinations of these skills through quantitative metrics, offering a nuanced perspective on a model’s proficiency in tackling intricate multimodal tasks.
TGIF-QA [20] adapts image question answering to the video domain, specifically targeting GIFs. With 165K question-answer pairs, it introduces tasks such as counting repetitions, identifying repeating actions, detecting state transitions, and frame-specific QA. These tasks demand detailed spatio-temporal analysis, making it a rigorous test for video comprehension. We employ GPT-3.5-turbo to assist in the evaluation of accuracy (same for the following three benchmarks). We denote it as “TGIF” in the main text.
MSVD-QA [51] is built on the Microsoft Research Video Description (MSVD) dataset [6], which includes 1,970 video clips and about 50,500 QA pairs. Its open-ended questions span five types—what, who, how, when, and where—offering a diverse and widely used evaluation for video question answering and captioning tasks. We denote it as “MSVD” in the main text.
MSRVTT-QA [6] comprises 10K video clips and 243K QA pairs, presenting a complex challenge due to the need to process both visual and temporal information. Like MSVD-QA, it features five question types, testing a model’s ability to understand and reason about dynamic video content. We denote it as “MSRVTT” in the main text.
ActivityNet-QA [58] offers 58K human-annotated question-answer pairs across 5.8K videos from the ActivityNet dataset. Its questions include motion, spatial relationships, and temporal dynamics, requiring long-term spatio-temporal reasoning and making it an popular benchmark for evaluating advanced video understanding abilities. We denote it as “ActivityNet” in the main text.